Find all points (x,y) where the functions f(x),g(x),h(x) have the same value:

f(x) = 2x-5 + 3, g(x) = 2x-5, h(x) = 8/x + 10


1
Expert's answer
2020-11-16T19:54:48-0500

Firstly, let us find all points where the functions "g(x)=2x-5" and "h(x)=\frac{8}{x}+10" have the same value. For this solve the following equation:


"2x-5=\frac{8}{x}+10"


which is equivalent to


"2x^2-5x=8+10x"


"2x^2-15x-8=0"


"2x^2-16x+x-8=0"


"2x(x-8)+x-8=0"


"(2x+1)(x-8)=0"


"x=8" or "x=-\frac{1}{2}"


Since "f(8)=2^3+3=11" and "g(8)=h(8)=11", at point "(8,11)" the functions "f(x),g(x),h(x)" have the same value.


Since "f(-\frac{1}{2})=2^{-\frac{1}{2}-5}+3>0" and "g(-\frac{1}{2})=-6<0", in this case the functions "f(x),g(x),h(x)" have different values.


Answer: "(8,11)"